# Augmenting the action space with conventions to improve multi-agent cooperation in Hanabi

### v1

@@ 4.3 Preliminary Results on Small Hanabi
Small Hanabi has a deck reduced to 20 cards, i.e. only having two suits, is limited to only two players, and the player hand size has been reduced to two cards. Furthermore, the life tokens have been reduced to one, and the hint tokens to three, otherwise the game shares all the core mechanics as regular Hanabi. This results in the problem having a significantly smaller state-action space and allows for initial testing and development of algorithms, with the aim of solving the full Hanabi problem. Even though this problem is considered simpler than Hanabi, it is by no means an easy problem to solve, with a perfect score of 10/10 being difficult to achieve. Player hands are often “locked” with important cards that must not be discarded (called critical cards), hint tokens run out almost immediately, and players are often forced into no-win scenarios where they must choose a bad action or risk losing their single life token.
In these tests, we use our in-house learning environment as well as the open-sourced learning environment developed by Bard et al. [36]. We implement our own rudimentary conventions specific to Small Hanabi, which we developed through multiple human plays. Fig. 3(a) shows the result for independent Deep Q-learning (DQN) with a primitive-action space compared to independent Deep Q-learning with a pure convention [^10] A pure convention space accounts for all possible scenarios, and the agents will always have a convention available. In Small Hanabi this is possible due to the small state-action space, but becomes a far greater challenge in the full Hanabi problem. space. DQN with a pure convention space performs significantly better than DQN with a primitive-action space, achieving a faster training time and improved converged performance. To validate these results, and improve on the performance further, we conduct an initial test using Rainbow and the open-sourced Hanabi learning environment with the Small Hanabi preset, the results are shown in Fig. 3(b). Rainbow is able to perform better than independent Deep Q-learning achieving a score of 7,6/10 compared to 5/10, and when substituting the primitive-action space for a pure convention space, the agents achieve significantly faster training time along with a slight convergent performance increase.
Figure 3: (a) Learning curves for independent Deep Q-learning (DQN) with a primitive-action space, pure conventions space, and an augmented and simplified primitive-action-convention space tested in our in-house Small Hanabi environment. (b) Learning curves for Rainbow with a primitive-action space, pure conventions space, and an augmented and simplified primitive-action-convention space tested in DeepMind’s Hanabi learning environment with the Small Hanabi preset [36].
@@
Due to Hanabi having a long list of intricate conventions, in addition to augmenting the action space with conventions, we experiment with the idea of simplifying the list of conventions during augmentation. This simplified list consists of the conventions that were the most straightforward and natural to develop, and were also the most common ones to occur in a game. The removed conventions can be considered edge-case or specific to certain game states that are very uncommon, and often results in uncertainty regarding their effectiveness. Fig. 3(a) and Fig. 3(b) shows the performance for DQN agents and Rainbow agents, respectively, with an augmented action-conventions-space. The performance impact is almost insignificant when compared to each method’s results for a pure convention space. This proves that even though only half of the conventions are present alongside primitive actions, the agents are still able to perform significantly better compared to only having primitive actions, and only slightly impacts performance compared to pure conventions. Therefore, when applying conventions in the full Hanabi problem, only a subset of established human conventions will be needed alongside primitive actions. Most existing Hanabi conventions have a list of core conventions which are built on the most essential principles, and are generally considered standardised. It is these conventions [^11] For an in depth discussion on each convention see the H-Group section I – VIII. alongside primitive actions that will be used when testing agents in the full Hanabi problem.
@@
It is important to note that there could exist a combination of conventions that would result in better performance uplifts, especially when a larger variety of conventions are implemented. However for the sake of this argument, we focused on the principle conventions since they are generally the least contentious and offer a good starting point for exploring the capabilities of action space augmentation with conventions in Hanabi. Further research can be conducted to explore more advance conventions, including multistep conventions, and their potential performance benefits.
@@ 5 Performance Evaluation Using Hanabi
The preliminary tests for action augmentation with conventions in Small Hanabi has shown promising results, and allow the agents to achieve faster training time and improved policies. We now shift focus to the full Hanabi problem for 2–5 players and with the principle conventions defined by the H-Group. Our results will focus on Bard et al.’s [36] Rainbow agent as a baseline, and apply an augmented action-convention space to improve on its performance. Rainbow was chosen as our baseline since it is among the methods that train the fastest, is the most sample efficient among the existing Hanabi agents [36, 13, 14, 40, 27], has little run to run variance [36], is considered a widely applicable algorithm [57, 58, 59], and still leaves considerable performance to be desired within Hanabi. All tests are conducted on the open-sourced Hanabi learning environment developed by Bard et al. [36].
@@ 5.1 Experimental Setup
Before discussing the results, we will briefly highlight the architecture design for each deep RL technique. For a full list of the hyperparameters used in each method see Table 4 in Appendix A. All methods receive a one-hot encoded observation defined by the environment as input to their neural networks, notably this observation tuple also contains the most recent actions within the previous round. The Rainbow agents use a Multilayer Perceptron (MLP) consisting of two feed forward neural networks with 512 neurons each, with the ReLU activation function [60] applied. It implements the Adam optimizer [54] to calculate the TD error and perform backpropagations to update the network weights. Additionally, the Rainbow agents use distributional reinforcement learning to predict the value distributions which are approximated as a discrete distribution over 51 uniform atoms, along with prioritise replay memory sampling [28]. Even though Rainbow uses n-step bootstrapping, in these experiments a value of n=1 was found to be optimal, additionally noisy nets have been disabled in favour for a traditional decaying epsilon-greedy approach.
When applying conventions to Rainbow, we only augment the action space and leave the reward signal and observation space untouched. Additionally, we add the necessary convention layers to the network architecture, but keep the core algorithm unaltered, as illustrated by Fig. 2. Through testing, we found the optimal hyperparameters to be similar to baseline Rainbow, as shown in Table 4 in Appendix A. During evaluation we compare each approach’s exponential moving averages (with a weight value of 0.9995) and standard error of the mean. All agents are trained using an Intel Core i7 10 700K CPU and Nvidia RTX 3080Ti GPU. The reward type for the Hanabi learning environment is set to non-lenient, i.e. the agents will receive a large negative reward (equal to the score) when bombing out.
@@ 5.2 Self-play Performance
The learning curves for Rainbow with and without an augmented action-convention space for 2–5 player Hanabi are shown in Fig. 4. The two player scenario seen in Fig. 4(a) shows the smallest improvement when applying conventions, however the two player Hanabi problem can be considered a special case. This is due to the fact that the shared information between the two players are far more limited when compared to higher player counts.
@@
For example, in two player Hanabi, half of the hands are hidden to a single player whereas in the three player scenario only a third of the hands are hidden, and the players always have a common hand visible. This allows for more advanced reasoning over another player’s actions, and can be seen in two of the principle conventions (specifically the Finesse and the Prompt) which cannot be applied in two player Hanabi.
Despite the fact that the complexity of the problem increases exponentially as more agents are added to the environment, the performance uplift of conventions become more apparent. In the three player setting seen in Fig. 4(b), the agents are able to achieve a significantly faster training time as well as an improved convergent performance with a higher average score. The performance improvement becomes even more apparent when looking at the five player setting, seen in Fig. 4(d), which is generally considered the most difficult problem to solve, and where the agents train roughly 10 times faster than the baseline Rainbow agents. In Fig. 5 the distribution of scores during evaluation of each agent over the course of 1 000 episodes is shown. In each scenario, conventions allow for a significant performance improvement, achieving a consistent grouping near a score of 21/25 and less overall variance when compared to baseline Rainbow. Rainbow with an augmented action-convention space also displays a robustness to increased player counts, with a remarkably consistent behaviour across all scenarios.
@@
Table 1 shows the average score for the best agent from each player count’s various runs compared to the results from Bard et al. [36]. Rainbow with an augmented action-convention space demonstrates an improved score for each scenario, with a higher average in the 3–5 player scenario and overall less deviation. Additionally, the four and five player results are competitive compared to other algorithms, such as ACHA, which requires 20 billion training steps to achieve a score of 16.8/25 in the five player setting.
@@
Upon investigation, the agents tend to choose conventions over primitive actions 70% of the time. This is the main reason for the Rainbow agents not achieving a higher average score, since they struggle to learn some of the core conventions available to them (specifically the Chop). We believe this to be a result of the Rainbow algorithm not having a memory of past observations, and therefore not being able to realise the importance of conventions that have a high payoff over extended time frames.
@@ 5.3 Cross-play Performance
Although self-play has historically led to the highest state-of-the-art agent performance within Hanabi [14], cross-play is also a crucial and widely applicable problem to solve. The ability for agents to cooperate across different training runs or regimes, or even across architectures, can greatly benefit MARL algorithms and their ability to cooperate in real world scenarios [47, 48]. Self-play agents are notorious for not being able to cooperate with never-before-seen partners, and often results in very poor performance [36, 27, 47, 48].
Table 2 shows the two player performance for various agents in cross-play Hanabi taken from existing literature, compared to the results of our cross-play Rainbow agents with an augmented action-convention space. Conventions are able to significantly improve on the performance of cross-play agents, allowing the baseline Rainbow agents to go from abysmal performance to an acceptable score. Furthermore, it is able to outperform SAD with only other-play applied for the two player scenario, with other-play and auxiliary task applied to SAD still holding the best score. However, as mentioned previously, auxiliary tasks only benefit the two player scenario in Hanabi [13]. When moving from self-play to cross-play, Rainbow with an augmented action-convention space only suffered a slight performance decrease and maintained remarkable consistency across various agent parings.
@@
Table 3 shows the results for cross-play Rainbow with an augmented action-convention space for all player counts of Hanabi. As seen in the self-play results for Rainbow conventions and confirmed in Table 3, the two player scenario is a special case, with the agents showing the biggest performance degradation when moving from self-play to cross-play, and overall worst score compared to the other cross-play results. This is again due to the more powerful conventions not being available in two player Hanabi, thus, it is important to look at all the player scenarios when evaluating cross-play performance.
@@ 6 Conclusion
Hanabi offers a complex and intricate problem for multi-agent reinforcement learning agents. It tests the ability to reason over another player’s actions and intensions, while maintaining a limited communication channel and partial observability. Existing MARL algorithms focus on complex architecture design and implement advanced algorithms capable of achieving remarkable performance. However, these algorithms are often computationally expensive, and require immense amounts of training data to learn convergent policies. In this paper, we focused on a different approach, exploring another core aspect of the Hanabi problem, namely conventions, and how to incorporate them into MARL. We showed how existing human conventions can be implemented in a MARL scenario using a special form of cooperative actions and action space augmentations.
Our main results show that conventions are able to significantly improve on the training time for Rainbow agents in Hanabi, and in the case of three to five players, also improve on the converged policy. Other Hanabi algorithms, such as SAD or MAPPO, require billions of training steps to reach convergent policies, whereas conventions reduced the number of training steps for Rainbow to below 30 million consistently. Additionally, conventions showed significant performance increase for cross-play agents, and allowed the Rainbow agents to go from not cooperating at all to being able to achieve decent performance across all player scenarios. For the most difficult scenario of five players, the cross-play Rainbow conventions agents trained on 50 million steps are able to outperform the ACHA self-play agents trained on 20 billion steps.
The biggest avenue for future research is exploring and developing convention discovery, similar to that of option discovery, which will allow agents to produce and define their own conventions as they train and learn. Our research serves as the basis for this future endeavour, proving the importance of conventions within MARL problems containing partial observability and limited communication.

### v2

@@ 4.3 Preliminary Results on Small Hanabi
To test the viability and capabilities of conventions, we conduct initial tests using the Small Hanabi environment, since it offers a smaller state-action space and allows for a pure convention space [^12] A pure convention space contains only conventions, i.e. no primitive actions, and accounts for all possible scenarios, ensuring the agents will always have a convention available. In Small Hanabi this is possible due to the small state-action space, but becomes a far greater challenge in the full Hanabi problem.. We compare a pure convention space and an augmented action-convention space, as discussed in Section 3.2, against their primitive action benchmarks. Due to Hanabi having a long list of intricate conventions, in addition to augmenting the action space with conventions, we experiment with the idea of simplifying the list of conventions during augmentation [^13] This idea stemmed from the discussion by Sutton et al. [19], where they explored the capabilities of combining primitive actions with simplified options, since the primitive actions can account for the simplification at the slight cost of performance.. This simplified list consists of the conventions that were the most straightforward and natural to develop, and were also the most common ones to occur in a game. The removed conventions can be considered edge-case or specific to certain game states that are very uncommon, and often results in uncertainty regarding their effectiveness.
Small Hanabi has a deck reduced to 20 cards with only two suits, is limited to only two players, and the player hand size has been reduced to two cards. Furthermore, the life tokens have been reduced to one, and the hint tokens to three, otherwise the game shares all the core mechanics as regular Hanabi. This results in the problem having a significantly smaller state-action space and allows for initial testing and development of algorithms, with the aim of solving the full Hanabi problem. Even though this problem is considered simpler than Hanabi, it is by no means an easy problem to solve, with a perfect score of 10/10 being difficult to achieve. Player hands are often “locked” with important cards that must not be discarded (called critical cards), hint tokens run out almost immediately, and players are often forced into no-win scenarios where they must choose a bad action or risk losing their single life token.
In these tests, we use our in-house learning environment as well as the open-sourced learning environment developed by Bard et al. [36]. We implement our own rudimentary conventions specific to Small Hanabi, which we developed through multiple human plays, presented and discussed in Appendix A.1. Fig. 3(a) shows the result for independent Deep Q-learning (DQN) with a primitive-action space compared to independent Deep Q-learning with a pure convention space. DQN with a pure convention space performs significantly better than DQN with a primitive-action space, achieving a faster training time and improved converged performance. To validate these results, and improve on the performance further, we conduct an initial test using Rainbow and the open-sourced Hanabi learning environment with the Small Hanabi preset, the results are shown in Fig. 3(b). Rainbow is able to perform better than independent Deep Q-learning achieving a score of 7,6/10 compared to 5/10, and when substituting the primitive-action space for a pure convention space, the agents achieve significantly faster training time along with a slight convergent performance increase.
Figure 3: (a) Learning curves for independent Deep Q-learning (DQN) with a primitive-action space, pure conventions space, and an augmented and simplified action-convention space tested in our in-house Small Hanabi environment. (b) Learning curves for Rainbow with a primitive-action space, pure conventions space, and an augmented and simplified action-convention space tested in DeepMind’s Hanabi learning environment with the Small Hanabi preset [36].
When augmenting the action space with conventions, we removed conventions 4-7 discussed in Table 4 found in Appendix A.1. Fig. 3(a) and Fig. 3(b) shows the performance for DQN agents and Rainbow agents, respectively, with an augmented action-conventions space. The performance impact is almost insignificant when compared to each method’s results for a pure convention space. This demonstrates that even though only half of the conventions are present alongside primitive actions, the agents are still able to perform significantly better compared to only having primitive actions, and only slightly impacts performance compared to pure conventions. Therefore, when applying conventions in the full Hanabi problem, only a subset of established human conventions will be needed alongside primitive actions. Most existing Hanabi conventions have a list of basic conventions which are built on the most fundamental principles, and are generally considered standardised within the community. It is these conventions, as presented and discussed in Appendix A.2, alongside primitive actions that will be used when testing agents in the full Hanabi problem.
@@
It is important to note that there could exist a combination of conventions that would result in better performance uplifts, especially when a larger variety of conventions are implemented. However for the sake of this argument, we focused on the basic conventions since they are generally the least contentious, less complex than more advanced strategies, and offer a good starting point for exploring the capabilities of action space augmentation with conventions in Hanabi. Further research can be conducted to explore more advance conventions, including advanced multistep conventions, and their potential performance benefits.
@@ 5 Performance Evaluation Using Hanabi
The preliminary tests for action augmentation with conventions in Small Hanabi has shown promising results, and allow the agents to achieve faster training time and improved policies. We now shift focus to the full Hanabi problem for 2–5 players and with the conventions defined by the H-Group and discussed in Appendix A.2. Our results will focus on Bard et al.’s [36] Rainbow agent as a baseline, and apply an augmented action-convention space to improve on its performance. Rainbow was chosen as our baseline since it is among the methods that train the fastest, is the most sample efficient among the existing Hanabi agents [36, 13, 14, 40, 27], has little run to run variance [36], is considered a widely applicable algorithm [57, 58, 59], and still leaves considerable performance to be desired within Hanabi. All tests are conducted on the open-sourced Hanabi learning environment developed by Bard et al. [36].
@@ 5.1 Experimental Setup
Before discussing the results, we will briefly highlight the architecture design for each deep RL technique. For a full list of the hyperparameters used in each method see Table 6 in Appendix B. All methods receive a one-hot encoded observation defined by the environment as input to their neural networks, notably (by default) this observation tuple contains the most recent actions within the previous round, and is not added as an additional feature. The Rainbow agents use a Multilayer Perceptron (MLP) consisting of two feed forward neural networks with 512 neurons each, with the ReLU activation function [60] applied. It implements the Adam optimizer [55] to calculate the TD error and perform backpropagations to update the network weights. Additionally, the Rainbow agents use distributional reinforcement learning to predict the value distributions which are approximated as a discrete distribution over 51 uniform atoms, along with prioritise replay memory sampling [28]. Even though Rainbow uses n-step bootstrapping, in these experiments a value of n=1 was found to be optimal, additionally noisy nets have been disabled in favour for a traditional decaying epsilon-greedy approach.
When applying conventions to Rainbow, we only augment the action space and leave the reward signal and observation space untouched. Additionally, we add the necessary convention layers to the network architecture, but keep the core algorithm unaltered, as illustrated by Fig. 1. Through testing, we found the optimal hyperparameters to be similar to baseline Rainbow, as shown in Table 6 in Appendix B. During evaluation, we compare each approach’s exponential moving averages (with a weight value of 0.9995) and standard error of the mean. All agents are trained using an Intel Core i7 10700K CPU and Nvidia RTX 3080Ti GPU. The reward type for the Hanabi learning environment is set to non-lenient, i.e. the agents will receive a large negative reward (equal to the score) when bombing out.
@@ 5.2 Self-play Performance
The learning curves for Rainbow with and without an augmented action-convention space for 2–5 player Hanabi are shown in Fig. 4. The two player scenario seen in Fig. 4(a) shows the smallest improvement when applying conventions, however the two player Hanabi problem can be considered a special case. This is due to the fact that the shared information between the two players are far more limited when compared to higher player counts. For example, in two player Hanabi, half of the hands are hidden to a single player whereas in the three player scenario only a third of the hands are hidden, and the players always have a common hand visible. This allows for more advanced reasoning over another player’s actions, evident by the fact that two of the conventions, specifically the Finesse and the Prompt as discussed in Appendix A.2, cannot be applied in two player Hanabi.
@@
Despite the fact that the complexity of the problem increases exponentially as more agents are added to the environment, the performance uplift of conventions become more apparent. In the three player setting seen in Fig. 4(b), the agents are able to achieve a significantly faster training time as well as an improved convergent performance with a higher average score. The performance improvement becomes even more apparent when looking at the five player setting, seen in Fig. 4(d), which is generally considered the most difficult problem to solve, and where the agents train roughly five times faster than the baseline Rainbow agents. In Fig. 5 the distribution of scores during evaluation of each agent over the course of 1000 episodes is shown. In each scenario, conventions allow for a significant performance improvement, achieving a consistent grouping near a score of 21/25 and less overall variance when compared to baseline Rainbow. Rainbow with an augmented action-convention space also displays a robustness to increased player counts, with a remarkably consistent behaviour across all scenarios.
@@
Table 1 shows the average score for the best agent from each player count’s various runs compared to the results from Bard et al. [36]. Rainbow with an augmented action-convention space demonstrates an improved score over baseline Rainbow for each scenario, with a higher average in the 3–5 player scenario and overall less deviation. Additionally, the three and five player results are competitive compared to other algorithms, such as ACHA, while also requiring substantially less training steps. In the case of five-player Hanabi, Rainbow with conventions outperform ACHA where ACHA required 20 billion training steps and Rainbow with conventions only required 30 million. A statistical significance test, conducted in Appendix C, shows that majority of the results are statistically different. However, the Rainbow with conventions agent’s scores are statistically similar to that of baseline Rainbow two players and ACHA three players, however conventions allow the Rainbow algorithm to reach these scores significantly faster (5x compared to baseline Rainbow and 1000x compared to ACHA). Furthermore, even though these two instances are statistically similar, the agents learn completely different policies, as seen by the cross-play performance evaluation in the next section.
@@
Upon investigation, the agents tend to choose conventions over primitive actions 70% of the time. This is the main reason for the Rainbow agents not achieving a higher average score, since they struggle to learn some of the conventions available to them (specifically The Chop as discussed in Appendix A.2). We believe this to be a result of the Rainbow algorithm not having a memory of past observations, and therefore not being able to realise the importance of conventions that have a high payoff over extended time frames.
@@ 5.3 Cross-play Performance
Although self-play has historically led to the highest state-of-the-art agent performance within Hanabi [14], cross-play is also a crucial and widely applicable problem to solve. The ability for agents to cooperate across different training runs or regimes, or even across architectures, can greatly benefit MARL algorithms and their ability to cooperate in real-world scenarios [48, 49]. Self-play agents are notorious for not being able to cooperate with never-before-seen partners, and often results in very poor performance [36, 27, 48, 49].
Table 2 shows the two player performance for various agents in cross-play Hanabi taken from existing literature, compared to the results of our cross-play Rainbow agents with an augmented action-convention space. The results for the Rainbow with conventions agents are statistically different from the other results reported in Table 2, as seen by Table 7 in Appendix C. Conventions are able to significantly improve on the performance of cross-play agents, allowing the baseline Rainbow agents to go from abysmal performance to an acceptable score. Furthermore, it is able to outperform SAD with only other-play applied for the two player scenario, with other-play and auxiliary task applied to SAD still holding the best score. However, as mentioned previously, auxiliary tasks only benefit the two player scenario in Hanabi [13]. When moving from self-play to cross-play, Rainbow with an augmented action-convention space only suffered a slight performance decrease and maintained remarkable consistency across various agent parings.
@@
Table 3 shows the results for cross-play Rainbow with an augmented action-convention space for all player counts of Hanabi. As seen in the self-play results for Rainbow with conventions and confirmed in Table 3, the two player scenario is a special case, with the agents showing the biggest performance degradation when moving from self-play to cross-play, and overall worst score compared to the other cross-play results. While most research efforts on cross-play Hanabi focus on the two player scenario, our results demonstrate that more powerful conventions are applicable in higher player counts, thus, it is important to look at all the player scenarios when evaluating cross-play performance.
@@ 6 Conclusion
Hanabi offers a complex and intricate problem for multi-agent reinforcement learning agents, since it incorporates various features of real-world problems, such as those found in the autonomous agents controlling vehicles problem. It tests an agent’s ability to reason over another player’s actions and intensions, while maintaining a limited communication channel and partial observability. Existing MARL algorithms focus on complex architecture design and implement advanced algorithms capable of achieving remarkable performance. However, these algorithms are often computationally expensive, and require immense amounts of training data to learn convergent policies. In this paper, we focused on a different approach, exploring another core aspect of the Hanabi problem, namely conventions, and how to incorporate them into MARL. We showed how existing human conventions can be implemented in a MARL scenario using a special form of cooperative actions and action space augmentations.
Our main results show that conventions are able to significantly improve on the training time for Rainbow agents in Hanabi, and in the case of three to five players, also improve on the converged policy. Other Hanabi algorithms, such as SAD or MAPPO, require billions of training steps to reach convergent policies, whereas conventions reduced the number of training steps for Rainbow to below 30 million consistently. Additionally, conventions showed significant performance increase for cross-play agents, and allowed the Rainbow agents to go from not cooperating at all to being able to achieve decent performance across all player scenarios. For the most difficult scenario of five players, the cross-play Rainbow conventions agents trained on 50 million steps are able to outperform the ACHA self-play agents trained on 20 billion steps. Our work, as presented in this paper, mainly focused on single-and two-step conventions in the Hanabi problem, with the goal of extending these concepts to other multi-agent scenarios (including simultaneous action spaces) containing multistep conventions, as well as the exploration of more advance conventions (and the application of those conventions) on more advanced algorithms within the Hanabi setting.
An additional potential benefit of conventions is that of encrypted communication, where a malicious user tries to intercept the ad-hoc agent’s communications, for example military communication or team communication in sports. Conventions can allow for an additional means of communicating encrypted sensitive data, even if the main communication channel is compromised, since implicit communication takes place through the observed actions, and can be explored in future research. However, we believe the biggest avenue for future research is exploring and developing convention discovery, similar to that of option discovery, which will allow agents to produce and define their own conventions as they train and learn. We believe that our research will serve as the foundation for this future endeavour, proving the importance of conventions within MARL problems containing partial observability and limited communication.